# FoilDiff: A Hybrid Transformer Backbone for Diffusion-based Modelling of 2D Airfoil Flow Fields

### v1

@@ 2.2 Non-Markovian Inference
The Markovian nature of the conventional diffusion process introduces certain limitations. The step-by-step denoising operation requires a large number of iterative steps to generate high-quality samples. To address these limitations, researchers have proposed alternative approaches that relax the Markovian assumption, leading to more efficient and expressive generative processes. Since the generative model approximates the reverse of the inference process, rethinking the inference process is key to reducing the number of iterations required during generation. Non-Markovian (see Figure 2) inference introduces dependencies across multiple time steps rather than restricting transitions to adjacent steps [30]. This innovation allows for faster sampling since fewer iterations are required to traverse the latent space while still capturing the complexity of the target distribution. This forward process is expressed as:
@@
where $q(\mathbf{x}_{T}|\mathbf{x}_{0})=\mathcal{N}(\sqrt{\alpha_{T}}\mathbf{x}_{0}(1-\alpha_{T})I)$, for all $t>1$. With the mean given by:
$q(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\sqrt{\alpha_{t-1}}\mathbf{x}_{0}+\sqrt{1-\alpha_{t}-1-\sigma_{t}^{2}}\frac{\mathbf{x}_{t}-\sqrt{\alpha_{t}}\mathbf{x}_{0}}{\sqrt{1-\alpha_{t}}},\sigma_{t}^{2}I),$ (2.12)
@@
$\mathbf{x}_{t-1}=\sqrt{\alpha_{t-1}}{\left(\frac{\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\ \mathbf{\epsilon}_{\theta}(\mathbf{x}_{t})}{\sqrt{\alpha_{t}}}\right)}+{\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\mathbf{\epsilon}_{\theta}(\mathbf{x}_{t})}+{\sigma_{t}\mathbf{\epsilon}}$ (2.13)
In this context, $\mathbf{\epsilon}\sim\mathcal{N}(0,I)$ represents standard Gaussian noise, which is independent of $\mathbf{x}_{t}$. The selection of $\sigma_{t}$ values dictates the nature of the generative process, enabling diverse configurations without necessitating the re-training of the model noise predictor $\mathbf{\epsilon}_{\theta}$. Specifically, when $\sigma_{t}=\sqrt{(1-\alpha_{t-1})/(1-\alpha_{t})}\sqrt{1-\alpha_{t}/\alpha_{t-1}}$ for all $t$, the forward process adheres to the Markovian property, and the generative process aligns with a DDPM. Conversely, when $\sigma_{t}=0$ for all $t>1$, the forward process becomes deterministic, relying solely on $\mathbf{x}_{t-1}$ and $\mathbf{x}_{0}$, with the exception of $t=1$. In this case, the coefficient associated with the random noise $\mathbf{\epsilon}_{t}$ reduces to zero, removing stochasticity from the generative process.
@@
This approach is formalized as the Denoising Diffusion Implicit Model (DDIM) [30], an implicit probabilistic model [37], generating samples deterministically by transitioning from latent variables $\mathbf{x}_{T}$ to $\mathbf{x}_{0}$. Despite departing from the traditional stochastic diffusion process, DDIM is trained using the same objective function as DDPM. As the model does not share parameters across time steps $t$, the $L_{1}$ objective [32] introduced in Equation 2.10 can serve as a surrogate for the variational objective. This substitution allows for consistent reuse of the noise predictor across all time steps, thereby simplifying the generative process while preserving predictive accuracy.
@@ 2.3 Noise Predictor in Diffusion Models
The noise predictor, defined in Equation 2.7, plays a crucial role in the reverse diffusion process. Diffusion models inherently model conditional distributions $p(\mathbf{x}_{t-1}|\mathbf{x}_{t})$, where conditioning enables external information including class labels, captions, or physical parameters to influence the generative process [38, 15, 39]. Lyu et al. [40] introduced a formulation where physical parameters are integrated as conditions $\mathbf{c}$ in the noise predictor, resulting in the modified loss function:
$L_{NN}(\theta)=\mathbb{E}_{\mathbf{x},\mathbf{\epsilon}\sim\mathcal{N}(0,I)}\left[\|\mathbf{\epsilon}-\mathbf{\epsilon}_{\theta}(\mathbf{x}_{t},\mathbf{c},t)\|^{2}\right].$ (2.14)
U-Net and Transformer-based neural networks have become the predominant architectures for noise prediction in diffusion models. Originally introduced for biomedical image segmentation [9], the U-Net architecture has since become a popular network architecture for noise prediction in diffusion-based models [25, 32], due to its ability to reconstruct fine-grained details. A wide range of conditioning mechanisms can be seamlessly integrated into the U-Net architecture, enabling the model to incorporate external information such as class labels, physical parameters, or auxiliary modalities [41, 42, 43]. The U-Net structure consists of a contracting path and an expanding path, connected via skip connections. The contracting path reduces spatial resolution while increasing feature depth, enabling hierarchical feature extraction. Conversely, the expansive path reconstructs the output back to its original resolution with skip connections to retain spatial details. This structure ensures a balance between global context understanding and fine-detail preservation, which is crucial for denoising accuracy [44].
Although convolutional backbones like U-Net remain widely used, transformer-based architectures have emerged as strong alternatives for noise prediction in diffusion models. Diffusion Transformers (DiT) demonstrate superior generalization over U-Net based approaches, especially in large-scale generative tasks [27]. Their core advantage lies in capturing long-range dependencies through self-attention, formulated as:
@@
where $Q,K,V$ represent query, key, and value matrices derived from the encoded conditioning information and input features, and $d_{v}$ is the dimension of the value matrix. The self-attention mechanism [45] dynamically weighs different regions of an input feature map, allowing transformers to model complex dependencies beyond local receptive fields.
Figure 3: Forward diffusion process on an airfoil flow field, showing progressive noise corruption from the original flow field at $t=0$ to a fully noisy state at step $t=1000$
Beyond standard conditioning, integrating attention mechanisms within the U-Net framework enhances diffusion models’ adaptability. Inspired by transformer architectures [45], attention gates in U-Net layers enable flexible multimodal conditioning, improving synthesis quality and domain-specific performance [46, 47, 6]. Regardless of the architectural choice, the efficiency of diffusion models heavily depends on sampling strategies. Optimizing the number of denoising steps remains crucial, as each step requires running the noise predictor multiple times.
@@ 3 Model Architecture and Configuration
The FoilDiff framework integrates a hybrid model architecture, structured dataset, and efficient inference procedures to predict airfoil flow fields across diverse aerodynamic regimes. The overall design as seen in Figure 4 leverages a latent transformer, with U-Net-style connected encoder-decoder paths, trained under the canonical denoising diffusion probabilistic modelling paradigm [32]. Our implementation explores two separate latent transformer architectures and an accelerated DDIM sampling strategy to reduce the number of inference steps required while preserving predictive fidelity [30].
The model is conditioned on encoded physical parameters, including airfoil geometry, Reynolds number, and angle of attack. These parameters are further reinjected into the latent transformer to enable improved generalization across a wide range of aerodynamic conditions. The combination of hybrid architectural design, domain-aware conditioning, and efficient inference enables FoilDiff to achieve improved accuracy and uncertainty estimation compared to existing U-Net-based baselines.
The following subsections detail each component of the model architecture, training pipeline, and data preparation process.
@@ 3.1 Encoder and Decoder
The encoder and decoder components of FoilDiff are designed with U-Net-style skip connections to support a latent transformer. The encoder comprises a sequence of convolutional downsampling blocks, each consisting of two convolutional layers with group normalization and GELU activation, followed by a $2\times 2$ strided convolution for spatial reduction. Positional encodings and a multi-headedself-attention mechanism are added at each resolution level to retain spatial context throughout the hierarchy.
The decoder mirrors the encoder structure, employing transposed convolutional blocks for spatial upsampling, interleaved with convolutional layers. To further enhance feature retention, skip connections are incorporated from the encoder to the decoder. These connections facilitate the transfer of spatial information between the contracting and expansive paths, mitigating information loss during downsampling. Mathematically, the residual mapping introduced by the skip connections is given by:
$f(\mathbf{y})=\mathbf{y}+H({\mathbf{y}}),$ (3.1)
where $\mathbf{y}$ is the input feature map, $H(\mathbf{y})$ represents the learned transformation (a series of convolutional layers), and the output $f(\mathbf{y})$ is the element-wise sum of the input and its transformed counterpart.
These skip connections improve gradient flow during training, preventing vanishing gradient issues and aiding in the reconstruction of fine-grained flow structures. Studies [44, 12] have shown that such architectural enhancements lead to improved convergence rates and more stable training in deep learning-based fluid dynamics simulations.
@@ 3.2 Latent Transformer
At the core of FoilDiff are latent vision transformer blocks, positioned at the bottleneck of the U-Net-like architecture. These enable better global context modelling, and deeper conditioning of the latent by allowing all spatial locations in the latent feature map to attend to one another and the conditions, addressing the limited receptive field of strictly convolutional networks.
Figure 5: Latent transformer architectures. (a) Standard DiT architecture illustrating the conventional diffusion transformer pipeline [27]; (b) U-ViT architecture showing modifications [48]. $Z\in\mathbb{R}^{B\times M}$ denotes the latent representation obtained at the final block of the encoder, where $B$ is the batch size, $M$ the embedding dimension, and $D$ is the total number of blocks
The latent representation, obtained after a series of downsampling operations in the encoder, is first reshaped into a sequence of flattened patches. Each patch is linearly projected into an embedding space and augmented with learnable positional encodings to retain spatial order. These embeddings are then passed through a stack of transformer encoder layers as seen in Figure 5, each composed of multi-head self-attention and feed-forward sublayers, interleaved with layer normalization and residual connections scaled, reshaped, and modulated by the physical conditions, ensuring that the generative process remains grounded in the specified physical scenario.
@@ 3.3 Training and Inference
The final stage of the model is the evaluation step, which during training is carried out by minimizing the loss defined in Equation 2.14 between the predicted noise and the noise added during the diffusion process. While the FoilDiff framework is compatible with alternatives, such as the $x$ and $v$ parametrisations [49], all architectures, including both the baseline U-Net variants and the proposed FoilDiff model, are trained using the canonical denoising diffusion process with the standard $\mathbf{\epsilon}$-parametrisation [32]. This decision is motivated by the need to isolate the effect of architectural modifications, ensure a fair comparison with existing airfoil flow field diffusion models, and preserve the intended focus and scope of this study.
During inference, the iterative nature of diffusion models creates a central concern regarding efficiency. We implement a DDIM sampler with strided time step selection strategy defining a reduced subset with elements $s$ from the original full set $\mathbf{x}_{1},\dots,\mathbf{x}_{T}$, where the selection is based on an increment $n$, in the sequence [30]:
@@
3.4 Dataset and Preprocessing
The dataset used for training consists of flow field simulations around airfoils across a range of Reynolds numbers and angles of attack [50]. Each sample comprises a structured grid representation of the flow field, including pressure and velocity vectors, alongside the corresponding airfoil geometry $(\omega)$, Reynolds number $(Re)$, and angle of attack $(\alpha)$, and the resulting CFD data. To ensure a compact and numerically stable representation, the Reynolds number and angle of attack are transformed into a parametric encoding using $(Re\cos(\alpha))$ and $(Re\sin(\alpha))$, respectively. These normalized physical parameters are embedded as additional channels alongside a mask of the airfoil geometry ($\omega$) in a three-channel tensor, which conditions the model [7].
@@ 3.4 Dataset and Preprocessing
and the remaining two channels corresponding to the normalized velocity components:
@@
where $p_{i}$ is the normalized pressure at location $i$, $p_{i,\infty}$ is the reference (freestream) pressure, $|\mathbf{u}_{f}|$ is the freestream velocity magnitude used for normalization, and $u_{x,i}$ and $u_{y,i}$ are the $x$- and $y$-components of the velocity at $i$. These are concatenated with the conditions, resulting in the six channel $32\times 32]$ tensor shown in Figure 6. These normalized representations enhance stability during training and ensure consistency across varying flow conditions. The dataset incorporates multiple simulations of identical input configurations to capture aleatoric uncertainty inherent in turbulent flow simulations [7]. For each Reynolds number in the dataset, we compute the pointwise mean and standard deviation from 20 simulations under identical input conditions. The resulting fields, shown in Figure 7, serve as a reference for both qualitative and quantitative evaluation of predictive uncertainty in our generative framework.
Figure 7: Mean fields and associated uncertainties computed from 20 samples of individual cases in the training and test dataset
4 Results and Discussion
To evaluate the predictive accuracy and computational efficiency of the proposed model, we evaluate test cases using the RAF30 airfoil geometry across a range of Reynolds numbers, as summarized in Table 1 with cases in regions of high and low uncertainty within and beyond the training range. Training and inference are performed on a single NVIDIA A100 GPU. The goal is to assess the model’s ability to generate high-fidelity flow field predictions under varying aerodynamic conditions and capture the epistemic uncertainty present. This is measured via the mean squared error (MSE) on both the mean prediction field $(MSE_{\mu})$ and predicted uncertainty field $(MSE_{\sigma})$.
Table 1: Dataset case configurations, including Reynolds number, subset assignment, and uncertainty category
Case Reynolds Number Subset Uncertainty Category
0 $0.5\times 10^{6}$ Test Low
1 $1.5\times 10^{6}$ Training
2 $2.5\times 10^{6}$ Test
3 $3.5\times 10^{6}$ Training
4 $4.5\times 10^{6}$ Test
5 $5.5\times 10^{6}$ Training High
6 $6.5\times 10^{6}$ Training
7 $7.5\times 10^{6}$ Test
8 $8.5\times 10^{6}$ Training
9 $9.5\times 10^{6}$ Test
10 $10.5\times 10^{6}$ Test

### v2

@@ 2.2 Non-Markovian Inference
The Markovian nature of the conventional diffusion process introduces certain limitations. The step-by-step denoising operation requires a large number of iterative steps to generate high-quality samples. To address these limitations alternative approaches that relax the Markovian assumption have been proposed, leading to more efficient and expressive generative processes [maleki2025diffusionmodelsmathematicalintroduction]. Non-Markovian inference (see Figure 2) introduces dependencies across multiple time steps rather than restricting transitions to adjacent steps [song2022denoisingdiffusionimplicitmodels]. This innovation allows for faster sampling since fewer iterations are required to traverse the latent space while still capturing the complexity of the target distribution. This forward process is expressed as:
@@
where $q(\mathbf{x}_{T}|\mathbf{x}_{0})=\mathcal{N}(\sqrt{\overline{\alpha}_{T}}\mathbf{x}_{0},\ (1-\overline{\alpha}_{T})I)$, for all $t>1$. With the mean given by:
$q(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\sqrt{\overline{\alpha}_{t-1}}\mathbf{x}_{0}+\sqrt{1-\overline{\alpha}_{t-1}-\sigma_{t}^{2}}\ \ .\ \frac{\mathbf{x}_{t}-\sqrt{\overline{\alpha}_{t}}\mathbf{x}_{0}}{\sqrt{1-\overline{\alpha}_{t}}},\sigma_{t}^{2}I),$ (2.7)
@@
$\mathbf{x}_{t-1}=\sqrt{\overline{\alpha}_{t-1}}{\left(\frac{\mathbf{x}_{t}-\sqrt{1-\overline{\alpha}_{t}}\ \boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t})}{\sqrt{\overline{\alpha}_{t}}}\right)}+{\sqrt{1-\overline{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t})}+{\sigma_{t}\boldsymbol{\epsilon}}$ (2.8)
In this context, $\boldsymbol{\epsilon}\sim\mathcal{N}(0,I)$ represents standard Gaussian noise, which is independent of $\mathbf{x}_{t}$. The selection of $\sigma_{t}$ values dictates the nature of the generative process, enabling diverse configurations without necessitating the re-training of the model noise predictor $\boldsymbol{\epsilon}_{\theta}$. This approach is formalized as the Denoising Diffusion Implicit Model (DDIM) [song2022denoisingdiffusionimplicitmodels], an implicit probabilistic model [mohamed2017learningimplicitgenerativemodels], generating samples deterministically by transitioning from latent variables $\mathbf{x}_{T}$ to $\mathbf{x}_{0}$.
@@ 2.3 Noise Predictor in Diffusion Models
The noise predictor, defined in Equation 2.3, plays a crucial role in the reverse diffusion process. Diffusion models inherently model conditional distributions $p(\mathbf{x}_{t-1}|\mathbf{x}_{t})$, where conditioning enables external information including class labels, captions, or physical parameters to influence the generative process [pan2022arbitrarystyleguidanceenhanced, dhariwal2021diffusionmodelsbeatgans, luo2022understandingdiffusionmodelsunified]. Lyu et al. [lyu2022a] introduced a formulation where physical parameters are integrated as conditions $\mathbf{c}$ in the noise predictor, resulting in the modified loss function:
$L_{NN}(\theta)=\mathbb{E}_{\mathbf{x},\boldsymbol{\epsilon}\sim\mathcal{N}(0,I)}\left[\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},\mathbf{c},t)\|^{2}\right].$ (2.9)
U-Net and Transformer-based neural networks have become the predominant architectures for noise prediction in diffusion models. Originally introduced for biomedical image segmentation [ronneberger2015unetconvolutionalnetworksbiomedical], the U-Net architecture has since become a popular network architecture for noise prediction in diffusion-based models [rombach2022highresolutionimagesynthesislatent, Ho2020], due to its ability to reconstruct fine-grained details. The U-Net structure consists of a contracting path and an expanding path, connected via skip connections. The contracting path reduces spatial resolution while increasing feature depth, enabling hierarchical feature extraction. Conversely, the expansive path reconstructs the output back to its original resolution with skip connections to retain spatial details. This structure ensures a balance between global context understanding and fine-detail preservation, which is crucial for denoising accuracy [prasad2023unravelingtemporaldynamicsunet]. A wide range of conditioning mechanisms can be seamlessly integrated into the U-Net architecture, enabling the model to incorporate external information such as class labels, physical parameters, or auxiliary modalities [yan2022diffusion, yang2024diffusionmodelscomprehensivesurvey, fu2024unveilconditionaldiffusionmodels].
Although convolutional backbones like U-Net remain widely used, transformer-based architectures have emerged as strong alternatives for noise prediction in diffusion models. Diffusion Transformers (DiT) demonstrate superior generative capacity over U-Net-based approaches, especially in large-scale generative tasks [peebles2023scalablediffusionmodelstransformers]. Their core advantage lies in capturing long-range dependencies through self-attention, formulated as:
@@
where $Q,K,V$ represent query, key, and value matrices derived from the encoded conditioning information and input features, and $d_{v}$ is the dimension of the value matrix. The self-attention mechanism [Vaswani2017] dynamically weighs different regions of an input feature map, allowing transformers to model complex dependencies beyond local receptive fields. Beyond standard conditioning, integrating attention mechanisms within the U-Net framework enhances diffusion models’ adaptability. Inspired by transformer architectures [Vaswani2017], attention gates in U-Net layers enable flexible multimodal conditioning, improving synthesis quality and domain-specific performance [pmlr-v139-radford21a, raffel2023exploringlimitstransferlearning, hu2024generativepredictionflowfield]. The noise predictor adopted in FoilDiff is designed to integrate the strengths of state-of-the-art architectures, thereby maximising predictive accuracy and generative performance.
@@ 3 Model Architecture and Configuration
The FoilDiff framework integrates a hybrid model architecture, structured dataset, and efficient inference procedures to predict airfoil flow fields across diverse aerodynamic regimes. The overall design, as seen in Figure 3 leverages a latent transformer, with U-Net-style connected encoder-decoder paths, trained under the canonical denoising diffusion probabilistic modelling paradigm [Ho2020]. Our implementation explores two separate latent transformer architectures and an accelerated DDIM sampling strategy to reduce the number of inference steps required while preserving predictive fidelity [song2022denoisingdiffusionimplicitmodels].
The model is conditioned on encoded physical parameters, including airfoil geometry, Reynolds number, and angle of attack. These parameters are further reinjected into the latent transformer to enable improved generalisation across a wide range of aerodynamic conditions. The combination of hybrid architectural design, domain-aware conditioning, and efficient inference enables FoilDiff to achieve improved accuracy and uncertainty estimation compared to existing U-Net-based baselines. The following subsections detail the model architecture and the methods employed during training and inference.
@@ 3.1 Encoder and Decoder
The encoder and decoder components of FoilDiff are designed with U-Net-style skip connections to support a latent transformer. The encoder comprises a sequence of convolutional downsampling blocks, each consisting of two convolutional layers with group normalization and GELU activation, followed by a $2\times 2$ strided convolution for spatial reduction. Positional encodings and a multi-head self-attention mechanism are added at each resolution level to retain spatial context throughout the hierarchy.
The decoder mirrors the encoder structure, employing transposed convolutional blocks for spatial upsampling, interleaved with convolutional layers. To further enhance feature retention, skip connections are incorporated from the encoder to the decoder. These connections facilitate the transfer of spatial information between the contracting and expansive paths, mitigating information loss during downsampling.
These skip connections improve gradient flow during training, preventing vanishing gradient issues and aiding in the reconstruction of fine-grained flow structures. Studies [prasad2023unravelingtemporaldynamicsunet, Wang2024] have shown that such architectural enhancements lead to improved convergence rates and more stable training in deep learning-based fluid dynamics simulations.
@@ 3.2 Latent Transformer
At the core of FoilDiff are latent vision transformer blocks, positioned at the bottleneck of the U-Net-like architecture. These enable better global context modelling and deeper conditioning of the latent by allowing all spatial locations in the latent feature map to attend to one another and the conditions, addressing the limited receptive field of strictly convolutional networks.
Figure 4: Latent transformer architectures. (a) Standard DiT architecture illustrating the conventional diffusion transformer pipeline [peebles2023scalablediffusionmodelstransformers]; (b) U-ViT architecture showing modifications [bao2023worthwordsvitbackbone]. $Z\in\mathbb{R}^{B\times M}$ denotes the latent representation obtained at the final block of the encoder, where $B$ is the batch size, $M$ the embedding dimension, and $D$ is the total number of blocks
The latent representation, obtained after a series of downsampling operations in the encoder, is first reshaped into a sequence of flattened patches. Each patch is linearly projected into an embedding space and augmented with learnable positional encodings to retain spatial order. These embeddings are then passed through a stack of transformer encoder layers as seen in Figure 4, each composed of multi-head self-attention and feed-forward sublayers, interleaved with layer normalization and residual connections scaled, reshaped, and modulated by the physical conditions using AdaLN-Zero [peebles2023scalablediffusionmodelstransformers]. This latent conditioning ensures that the generative process remains grounded in the specified physical scenario.
@@ 3.3 Training and Inference
The final stage of the model is the evaluation step, which during training is carried out by minimizing the loss defined in Equation 2.9 between the predicted noise and the noise added during the diffusion process. While the FoilDiff framework is compatible with alternatives, such as the $x$ and $v$ parametrisations [salimans2022progressivedistillationfastsampling], all architectures, including both the baseline U-Net variants and the proposed FoilDiff model, are trained using the canonical denoising diffusion process with the standard $\boldsymbol{\epsilon}$-parametrisation [Ho2020]. This decision is motivated by the need to isolate the effect of architectural modifications, ensure a fair comparison with existing airfoil flow field diffusion models, and preserve the intended focus and scope of this study.
During inference, the iterative nature of diffusion models creates a central concern regarding efficiency. We implement a DDIM sampler with strided time step selection strategy defining a reduced subset with elements $s$ from the original full set $\mathbf{x}_{1},\dots,\mathbf{x}_{T}$, where the selection is based on an increment $n$, in the sequence [song2022denoisingdiffusionimplicitmodels]:
@@
4 Experimental Analysis
This section presents a comprehensive evaluation of FoilDiff, focusing on its predictive accuracy, uncertainty quantification, and computational efficiency in comparison with existing state-of-the-art diffusion-based models. For training and evaluation, we utilize a dataset consisting of flow field simulations across a range of Reynolds numbers and angles of attack from a publicly available repository [TUM-IRepository1]. Each sample in this dataset comprises a structured grid representation of the flow field, including pressure and velocity vectors, alongside the corresponding airfoil geometry $(\omega)$, Reynolds number $(Re)$, and angle of attack $(\phi)$, and the resulting CFD data. To ensure a compact and numerically stable representation, the Reynolds number and angle of attack were transformed into a parametric encoding using $(Re\cos(\phi))$ and $(Re\sin(\phi))$, respectively. These normalised physical parameters were embedded as additional channels alongside a mask of the airfoil geometry ($\omega$) in a three-channel tensor, which conditions the model [Liu2023].
@@ 4 Experimental Analysis
and the remaining two channels corresponding to the normalised velocity components:
@@
where $p_{i}$ is the normalised pressure at location $i$, $p_{i,\infty}$ is the reference (freestream) pressure, $|\mathbf{u}_{f}|$ is the freestream velocity magnitude used for normalization, and $u_{x,i}$ and $u_{y,i}$ are the $x$- and $y$-components of the velocity at $i$. These were concatenated with the conditions, resulting in the six channel $128\times 128]$ tensor shown in Figure 5 which is easily downsampled to $64\times 64]$ and $32\times 32]$ for this study. These normalised representations enhance stability during training and ensure consistency across varying flow conditions.
Our experiments are designed to provide a fair and systematic benchmark by following evaluation protocols consistent with prior studies in diffusion-based flow field modelling. We aim to assess both the quantitative performance and the physical fidelity of FoilDiff’s predictions across a range of aerodynamic conditions.